1. Apples are collected in a basket for six people. One third, one fourth, one eight and one fifth are given to four people respectively. The fifth person gets ten apples with one apple remaining for the sixth person. Find the original number of apples in the basket?
Thank you very much for the answer ! :)

Answer by mananth(16949)

let number of apples be x
x/3 + x/4 + x/5 +x/8 + 11= x
LCD=480
multiply by LCD
160x+120x+96x+60x+5280=480x
436x-480x=-5280
-44x=-5280
x=120 apples